Preparing to switch to new storage blocks.
Restarting storage. Implementing alternative storage solutions.
Investigating storage slowdown.
Normal Service restored.
Services ready, will begin routing traffic back to pool momentarily.
Network protocol to the primary storage block is experiencing issues.
Team is working on restoring service.
Unresolved Sites using timthumb are still being investigated for performance issues.
Sites using timthumb are experiencing slowdowns, we are investigating.
Traffic coming thru our UK nodes are properly resolving to pool10.
A handful of requests routed through our UK facility are not resolving properly to the new pool. Correcting now.
Investigating a handful of site not resolving correctly.
Resuming normal operations.
Routing traffic here temporarily while we make a change to the db master configs.
Database cluster has stabilized. All systems operating normally and well within spec.
Adding additional database capacity.
Some database slowdown on the slaves, Engineers working on it.
Traffic is now resolving to pool10, please contact support with any issues. We are correcting a few anomalies as we see them.
Running final checks before switching traffic.
We are delaying slightly before switching the traffic over.
Engineers will be making the sync of the databases and final system prior to redirecting all traffic.
Earlier today we sent the following email to approx 200 customers. This blog post will track the status of the migration.